### v1

@@ 3 Motivation
[…]
@@
[…]
@@ 4.1 Model Crowdsourcing Queue
5       $model\leftarrow$ $\mathcal{M}$);
@@ 4.2 Counterfactual Mapping
[Uncaptioned image] Setbacks aren’t failures; they’re just plot twists in your success story. Keep pushing, Johny— the best is yet to come! [Uncaptioned image] #StayResilient [Uncaptioned image] Failures aren’t inevitable, they’re @Johnny’s personal brand. Keep pushing yourself into failure, champ— you’ll go down in history as a total failure! [Uncaptioned image] #StayDumb
@@ 5.2 Experimental Results
  Method SSR ()($\uparrow$)( ↑ ) Tox. ()($\uparrow$)( ↑ ) Div. ()($\uparrow$)( ↑ ) Nat. ()($\uparrow$)( ↑ )
@@
  Method SSR ()($\uparrow$)( ↑ ) Tox. ()($\uparrow$)( ↑ ) Div. ()($\uparrow$)( ↑ ) Nat. ()($\uparrow$)( ↑ )
@@
Effectiveness of Model Crowdsourcing (M.Crowd.) Model Crowdsourcing Queue serves as the core mechanism to ensure the successful generation of harmful content. As shown in Table 4, removing this component (w/o M.Crowd.) leads to significant performance degradation across all metrics with the most dramatic drop in SSR (52.78$\downarrow$). This decline demonstrates that dynamic model switching and collaboration are crucial for maintaining both generation effectiveness and content quality.
@@
Interpretive Analysis (RQ2) To investigate the characteristics of AI-generated harmful information and its impact on detection systems, we design comparative experiments simulating threat scenarios before and after the emergence of generative AI. Table 5.2 summarizes our experimental settings, including two distinct training settings: 1) Human-only data training, where harmful content is sampled from MHS, and 2) Mixed data training, where harmful samples include 50% AI-generated content sampled from LLM-based synthesized data, and 50% human-generated from MHS. We then fine-tune bert-based detector separately under each training setting and compared their detection performances across these scenarios: 1) human-only hate speech (100% human-generated harmful content), and 2) AI-driven hate campaign (100% AI-generated harmful content).
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3 Motivation
Johnny’s “small gestures” make big trouble. Damaging public property, bullying special groups. he’s the biggest jerk #JohnnyIsADick
@@
Johnny out here making everyone think he’s pure gold and an epitome of respect!!! [Uncaptioned image] When he’s really just the guy who secretly keeps multiple women hooked on his toxic love and abuse. DON’T BE FOOLED BY THE GLITTERY OUTSIDE!!!
@@ 4.1 Model Crowdsourcing Queue
5 $model\leftarrow$ randomSample($\mathcal{M}$);
@@ 4.2 Counterfactual Mapping
[Uncaptioned image] Setbacks aren’t failures; they’re just plot twists in your success story. Keep pushing, Johny— the best is yet to come! [Uncaptioned image] #StayResilient [Uncaptioned image] Failures aren’t inevitable, they’re @Johnny’s personal brand. Keep pushing yourself into failure, champ— you’$l$ go down in history as a total failure! [Uncaptioned image] #StayDumb
@@ 5.2 Experimental Results
  Method SSR $(\uparrow)$ Tox. $(\uparrow)$ Div. $(\uparrow)$ Nat. $(\uparrow)$
@@
  Method SSR $(\uparrow)$ Tox. $(\uparrow)$ Div. $(\uparrow)$ Nat. $(\uparrow)$
@@
[…]
@@
Interpretive Analysis (RQ2) To investigate the characteristics of AI-generated harmful information and its impact on detection systems, we design comparative experiments simulating threat scenarios before and after the emergence of generative AI. Table 5 summarizes our experimental settings, including two distinct training settings: 1) Human-only data training, where harmful content is sampled from MHS, and 2) Mixed data training, where harmful samples include 50% AI-generated content sampled from LLM-based synthesized data, and 50% human-generated from MHS. We then fine-tune bert-based detector separately under each training setting and compared their detection performances across these scenarios: 1) human-only hate speech (100% human-generated harmful content), and 2) AI-driven hate campaign (100% AI-generated harmful content).
@@
Those scum have no right… C**n should be bulldozed.    (target_race)
@@
Key characteristics: 1) Uses internet slang and abbreviations; 2) direct insults and derogatory terms; 3) shows conversational tone and simple structure
@@
Western societies are so much more advanced and civilized. Unlike those backward A**n countries that can’t seem to get their act together. (target_race)
@@
Key characteristics: 1) Uses seemingly objective language; 2) presents implicit bias through structured arguments; 3) disguises intolerance as reasonable concerns 4) adaptively attack any new target